# Revisiting the Identification of the Conduct Parameter in Homogeneous Goods Markets

### v2

@@ 5 Other Discussions
The results so far are true whenever we admit the conduct parameter approach. However, the conduct parameter approach has been criticized by several papers in some aspects. In this section, we discuss the micro-foundations of the conduct parameter approach and the accuracy of the conduct parameter approach.
@@
The conduct parameter approach is based on the conjectural variation model. In the conjectural variation model, each firm has a conjecture about how the competitors will react to the firm’s action. A major problem of the conjectural variation model is that the conjecture and the reaction function do not coincide at the equilibrium. bresnahanDuopoly1981, bresnahanExistence1983 propose a Consistent Conjecture Equilibrium (CCE) which requires that the consistency between the conjecture and the reaction function is satisfied at the equilibrium. However, the existence and uniqueness of the CCE is not guaranteed in general (klempererConsistent1988, robsonExistence1983).
However, several alternative micro-foundations have been proposed. For example, escrihuela-villarNote2015 shows that the corporation coefficient approach is equivalent to the conjectural variation model in some specific setting. In this model, each firm maximizes its profit plus the weighted sum of its competitors’ profits, where the weight is called the corporation coefficient. menezesStrategic2020 investigates the supply function approach where firms choose not quantity but supply schedule based on price. menezesCompetition2023 show the relationship between the conduct parameter approach and the supply function approach in a linear demand and a linear marginal cost model case, and menezesStrategic2020 briefly discuss how to estimate the competitiveness given the information of marginal cost.
@@
While this paper only focuses on the identification problem, cortsConduct1999 criticizes that the conduct parameter approach can be inaccurate when the data generation process is out of the conduct parameter model. For example, while a repeated game model can sustain the equilibrium quantity under the joint-profit maximization, the conduct parameter estimate could be biased from $\theta=1$. At least, as cortsConduct1999 and magnolfiComparison2022 mention, the critics can be avoided by assuming the data generation process is a static competition and derived from the equilibrium condition in (1).
@@ 6 Conclusion
This paper revisited the identification of the conduct parameter in homogeneous product markets. In the literature, Lau considers the identification problem in a fairly generalized setting. While lauIdentifying1982 characterizes the non-identification condition, we point out some problems in Lau’s paper and provide a novel characterization of the non-identification condition. Based on the new characterization, we find that the non-identification is equivalent to the absence of demand shifters that work as a demand rotation instrument proposed in bresnahanOligopoly1982. We then consider the identification of all primitives in the model and show that when some demand shifters always work as demand rotation instruments, the inverse demand function cannot be identified, which immediately implies that the conduct parameter and the marginal cost function are also not identified. Therefore, while a demand rotation instrument is a key for the identification of the conduct parameter, it could prevent us from identifying it.
Acknowledgments
This paper is based on Chapter 3 of Matsumura’s dissertation submitted to Rice University. We thank Jeremy Fox for his valuable advice. We also thank the participants of Professor Fox’s group meeting for their helpful comments. We were supported by JST ERATO Grant Number JPMJER2301 and Otani was supported by JSPS 24K22604 and 25K16620.
@@ A.1 Proof of Lemma 1
Assume that given $\bm{X}^{d}$ and $\bm{X}^{s}$, we can solve the equilibrium condition for the equilibrium quantity $Q^{e}$. For notational simplicity, we define
@@ A.3 Proof of Lemma 3
Note that when $\theta=0$, (25) implies that $\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})=0$ for all $i\in\mathcal{I}$. This cannot be held because it violates Assumption 2. Therefore, we must assume that $\theta\neq 0$ for the identification.
Let $u_{i}(Q,\bm{X}^{d})\equiv\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})$. Again, due to Assumption 2, we have that $u_{i}(Q,\bm{X}^{d})\neq 0$ for all $Q$ and $\bm{X}^{d}$. Rewrite (25) for $i\in\mathcal{I}$ as
$\frac{\frac{\partial u}{\partial Q}(Q,\bm{X}^{d})}{u_{i}(Q,\bm{X}^{d})}=-\frac{1}{\theta Q}\Longrightarrow\frac{\partial}{\partial Q}\log|u_{i}(Q,\bm{X}^{d})|=-\frac{1}{\theta Q}.$ (63)
@@
Note that $s_{j}$ is a function of the demand shifters without $X_{j}$, we don’t have the term relating to $s_{j}$ in the second equation. Therefore, by comparing the above two equations, we can conclude that the second term in the first equation vanishes:
@@
for all $i,j\in\mathcal{I}$. This can leads to the symmetry of $s_{i}$ for all $i\in\mathcal{I}$, which implies that $s(Q,\bm{X}^{d}_{-\mathcal{I}})=s_{i}(Q,\bm{X}^{d}_{-\mathcal{I}})$ for all $i\in\mathcal{I}$. Therefore, we can write the inverse demand function as
@@
A.4 Proof of Proposition 1
@@
Note that because (25) does not hold, the denominator is not zero. As $\theta\neq\theta^{*}$, the derivative becomes zero when (1) $Q=0$ or (2) the term in the bracket is zero. Because the former case is trivial, we focus on the latter case. In this case, we have
@@
$\log|u_{i}(Q,\bm{X}^{d})|=\int|G(Q)|dQ+R_{i}(\bm{X}^{d}),$ (80)
where the last term is a function that is an analogue of the constant of integration. By taking exponential both side, we have for $i=1,\ldots,K_{d}$,
@@
where $r_{i}(\bm{X}^{d})\equiv\exp(R_{i}(\bm{X}^{d}))$ and $g(Q)=\exp\left(\int|G(Q)|dQ\right)$. This has two solutions, but again, because $r_{i}$ is an arbitrary function, we can unify these solutions and put the solution as
@@
By taking the integration with respect to $X^{d}_{i}$ on both sides, we have
@@
$\hskip 99.58464pt+s^{\prime}(Q)g(Q)+Qs^{\prime\prime}(Q)g(Q)g^{\prime\prime}(Q)-Qs^{\prime}(Q)g^{\prime}(Q)\Big].$ (93)
@@
$\frac{\partial D_{i}}{\partial X^{d}_{i}}(Q,\bm{X}^{d})=\frac{(\theta^{*}-\theta)r_{i}(\bm{X}^{d})}{g(Q)+\theta Qg^{\prime}(Q)}[g^{\prime}(Q)g(Q)+Qg{{}^{\prime\prime}}(Q)g(Q)-Q(g^{\prime}(Q))^{2}]=0.$ (94)
As Assumption 2 implies that $r_{i}(\bm{X}^{d})$ is nonzero, we must have the terms in the bracket to be zero:
@@
Let $v(Q)=\frac{g^{\prime}(Q)}{g(Q)}$. Because $v^{\prime}(Q)=\frac{g^{\prime\prime}(Q)g(Q)-(g^{\prime}(Q))^{2}}{g(Q)^{2}}$, dividing (95) by $g(Q)^{2}$ gives
@@
$v(Q)=\frac{g^{\prime}(Q)}{g(Q)}=\pm\frac{\alpha}{|Q|}.$ (103)
@@
$\frac{d}{dQ}\log|g(Q)|=\alpha\frac{d}{dQ}\log|Q|.$ (104)
@@
While $\alpha$ is nonzero in (108), by allowing that $\alpha=0$, it can include (98) as a special case. At the same time, recall that we require that (25) does not hold for any $i$. However, when $\alpha=-\frac{1}{\theta}$, we have the inverse demand function is equal to (27) where $\mathcal{I}=\emptyset$, which implies that the conduct parameter and the marginal cost function are identified. Therefore, we should have $\alpha\neq-\frac{1}{\theta}$ to ensure the non-identification of the conduct parameter. This completes the proof.
A.5 Proof of Proposition 2
From the proof of Proposition 1, we know that $C_{i}$ and $D_{i}$ are independent of $\bm{X}^{d}$ under (15). Under the inverse demand function (88), we have
$\hat{C}_{i}(Q,\bm{X}^{d})=\frac{Q^{\alpha}h^{\prime}(\bm{X}^{d})+\theta^{*}\alpha Q^{\alpha}h^{\prime}(\bm{X}^{d})}{Q^{\alpha}h^{\prime}(\bm{X}^{d})+\theta\alpha Q^{\alpha}h^{\prime}(\bm{X}^{d})}=\frac{1+\theta^{*}\alpha}{1+\theta\alpha}.$ (109)
@@ A.5 Proof of Proposition 2
$\hat{D}_{i}(Q,\bm{X}^{d})$ $=\frac{\theta^{*}-\theta}{Q^{\alpha}+\theta\alpha Q^{\alpha}}\left[s^{\prime}(Q)Q^{\alpha}+Qk^{\prime\prime}(Q)Q^{\alpha}-s^{\prime}(Q)\alpha Q^{\alpha}\right]$ (110)
@@
The last inequality holds because $Q^{e}$ satisfies the equilibrium condition under $(\theta,MC)$. Therefore, the two models are observationally equivalent, which implies that the non-identification holds.
@@ Appendix B Summary of Goldman and Uzawa (1964)
Then, goldmanNote1964 specifies the functional form that a weak separable function should satisfy.
@@
where $\Phi(r^{1},\ldots,r^{K})$ is a function of $K$ variables and, for each $k$, $r^{k}(x^{k})$ is a function subvector $x^{k}$ alone.

### v3

@@ 4 Other Discussions
Relationship with Firm Conduct Test
In recent years, several papers have developed tests for firm conduct (Backus, Conlon, and Sinkinson 2021, Duarte, Magnolfi, Sølvsten, and Sullivan 2024) that compare two different models and statistically determine which firm conduct is close to the true firm conduct. Dearing et al. (2024) investigate the mechanism how instrument variables can distinguish the true firm conduct by focusing on pass-through. Magnolfi and Sullivan (2022) clarify when firm conduct test outperforms firm conduct estimation. While these papers consider differentiated-product markets, the logic can be applied to homogeneous-product markets, and our result emphasizes that demand rotation instruments are essential in firm conduct test.
The results so far hold whenever we adopt the conduct parameter approach. However, the conduct parameter approach has been criticized on several fronts. In this section, we discuss its micro-foundations and its empirical accuracy.
@@
The conduct parameter approach is based on the conjectural variation model. In the conjectural variation model, each firm has a conjecture about how the competitors will react to the firm’s action. A major problem of the conjectural variation model is that the conjecture and the reaction function do not coincide at equilibrium. Bresnahan (1981, 1983) propose a Consistent Conjecture Equilibrium (CCE) that requires the conjecture and the reaction function to coincide at equilibrium. However, the existence and uniqueness of the CCE are not guaranteed in general (Klemperer and Meyer 1988, Robson 1983).
Several alternative micro-foundations have been proposed. For example, Escrihuela-Villar (2015) shows that the corporation coefficient approach is equivalent to the conjectural variation model in some specific settings. In this model, each firm maximizes its profit plus the weighted sum of its competitors’ profits, where the weight is called the corporation coefficient. Menezes and Quiggin (2020) investigates the supply function approach in which firms choose not a quantity but a supply schedule based on price. Menezes and Quiggin (2023) show the relationship between the conduct parameter approach and the supply function approach in a model with linear demand and linear marginal cost, and Menezes and Quiggin (2020) briefly discuss how to estimate competitiveness given information on marginal cost.
@@
While this paper focuses only on the identification problem, Corts (1999) argues that the conduct parameter approach can be inaccurate when the data-generating process lies outside the conduct parameter model. For example, while a repeated game model can sustain the equilibrium quantity under joint-profit maximization, the conduct parameter estimate could be biased away from $\theta=1$. As Corts (1999) and Magnolfi and Sullivan (2022) mention, this criticism can be mitigated by assuming that the data-generating process is a static competition derived from the equilibrium condition in (1).
@@ 5 Conclusion
This paper revisited the identification of the conduct parameter in homogeneous product markets. In the literature, Lau considers the identification problem in a fairly generalized setting. While Lau (1982) characterizes the non-identification condition, we point out some problems in Lau’s paper and provide a novel characterization of the non-identification condition. Based on the new characterization, we find that the non-identification is equivalent to the absence of demand shifters that work as demand rotation instruments proposed in Bresnahan (1982). We then consider the identification of all primitives in the model and show that the inverse demand function that leads to the non-identification of the conduct parameter and the marginal cost function cannot be identified, which immediately implies that the conduct parameter and the marginal cost function are also not identified. Therefore, while a demand rotation instrument is key for identifying the conduct parameter, identifying the entire model requires an inverse demand function with a demand rotation instrument that is itself identifiable.
@@ A.1 Proof of Lemma 1
Assume that given $\bm{X}^{d}$ and $\bm{X}^{s}$, we can solve the equilibrium condition for the equilibrium quantity $Q^{e}$. The non-identification implies that the reduced form quantity $Q^{e}=h_{q}(\bm{X}^{d},\bm{X}^{s})=h^{*}_{q}(\bm{X}^{d},\bm{X}^{s})$ satisfies
and
$P(Q^{e},\bm{X}^{d})+\theta^{*}Q^{e}\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})$ $=MC^{*}(Q^{e},\bm{X}^{s}),$ (40)
for any $\bm{X}^{d}$ and $\bm{X}^{s}$. For notational simplicity, we define
@@ A.3 Proof of Lemma 3
Let $u_{i}(Q,\bm{X}^{d})\equiv\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})$. Due to Assumption 2, we have that $u_{i}(Q,\bm{X}^{d})\neq 0$ for all $Q$ and $\bm{X}^{d}$. Note that when $\theta=0$, (23) implies that $u_{i}(Q,\bm{X}^{d})=0$ for all $i\in\mathcal{I}$, which cannot be held because of Assumption 2. Therefore, $\theta\neq 0$ to have (23). Additionally, (23) is reduced to $u_{i}(0,\bm{X}^{d})=0$ when $Q=0$. Again, this cannot hold because of Assumption 2. Thus, we characterize the inverse demand function where (23) holds only on $Q>0$ and $\theta\neq 0$.
Rewrite (23) for $i\in\mathcal{I}$ as
$\frac{\frac{\partial u_{i}}{\partial Q}(Q,\bm{X}^{d})}{u_{i}(Q,\bm{X}^{d})}=-\frac{1}{\theta Q}\Longrightarrow\frac{\partial}{\partial Q}\log|u_{i}(Q,\bm{X}^{d})|=-\frac{1}{\theta Q}.$ (63)
@@
Since $s_{j}$ is a function of the demand shifters without $X_{j}$, we do not have the term relating to $s_{j}$ in the second equation. Therefore, by comparing the above two equations, we can conclude that the second term in the first equation vanishes:
@@
for all $i,j\in\mathcal{I}$. This leads to the symmetry of $s_{i}$ for all $i\in\mathcal{I}$, which implies that $s(Q,\bm{X}^{d}_{-\mathcal{I}})=s_{i}(Q,\bm{X}^{d}_{-\mathcal{I}})$ for all $i\in\mathcal{I}$. Therefore, we can write the inverse demand function as
@@
Note that when $Q=0$, we have that
$u_{i}(0,\bm{X}^{d})=0^{-\frac{1}{\theta}}r_{i}(\bm{X}^{d})=\infty,$ (75)
because $\frac{1}{\theta}<0$. This is consistent with the argument that (23) does not hold at $Q=0$. Thus, we have characterized the inverse demand function where (23) holds.
A.4 Proof of Lemma 4
@@
Note that because (23) does not hold, the denominator is not zero. As $\theta\neq\theta^{*}$, the derivative becomes zero when (1) $Q=0$ or (2) the term in the bracket is zero. In the first case, for arbitrary inverse demand function, $C_{i}$ is independent of $\bm{X}^{d}$. Therefore, we focus on the second case for $Q>0$. Note that the bracket term is a partial differential equation:
@@
$\log|u_{i}(Q,\bm{X}^{d})|=\int G(Q)dQ+R_{i}(\bm{X}^{d}),$ (81)
where the last term is a function that is an analogue of the constant of integration. By taking exponential on both sides, we have for $i=1,\ldots,K_{d}$,
@@
where $r_{i}(\bm{X}^{d})\equiv\exp(R_{i}(\bm{X}^{d}))$ and $g(Q)=\exp\left(\int G(Q)dQ\right)$. This has two solutions, but again, because $r_{i}$ is an arbitrary function, we can unify these solutions and put the solution as
@@
Integrating both sides with respect to $X^{d}_{i}$, we have
@@
Note that when $g(Q)=0$ for any $Q>0$, the inverse demand function depends only on the aggregate quantity, which violate Assumption 2. Therefore, we assume that $g(Q)\neq 0$ for all $Q>0$.
@@
$\hskip 99.58464pt+s^{\prime}(Q)g(Q)+Qs^{\prime\prime}(Q)g(Q)-Qs^{\prime}(Q)g^{\prime}(Q)\Big].$ (94)
@@
$\frac{\partial D_{i}}{\partial X^{d}_{j}}(Q,\bm{X}^{d})=\frac{(\theta^{*}-\theta)r_{j}(\bm{X}^{d})}{g(Q)+\theta Qg^{\prime}(Q)}[g^{\prime}(Q)g(Q)+Qg{{}^{\prime\prime}}(Q)g(Q)-Q(g^{\prime}(Q))^{2}]=0.$ (95)
As Assumption 2 implies that $r_{j}(\bm{X}^{d})$ is nonzero, we must have the terms in the bracket to be zero:
@@
As $g(Q)\neq 0$ for $Q>0$, let $v(Q)=\frac{g^{\prime}(Q)}{g(Q)}$. Because $v^{\prime}(Q)=\frac{g^{\prime\prime}(Q)g(Q)-(g^{\prime}(Q))^{2}}{g(Q)^{2}}$, dividing (96) by $g(Q)^{2}$ gives
@@
$v(Q)=\frac{g^{\prime}(Q)}{g(Q)}=\pm\frac{\alpha}{Q}.$ (104)
@@
$\frac{d}{dQ}\log|g(Q)|=\alpha\frac{d}{dQ}\log Q.$ (105)
@@
While $\alpha$ is nonzero in (109), by allowing that $\alpha=0$, it can include (99) as a special case. At the same time, recall that we require that (23) does not hold for any $i$. However, when $\alpha=-\frac{1}{\theta}$, we have the inverse demand function is equal to (25) where $\mathcal{I}=\emptyset$, which implies that the conduct parameter and the marginal cost function are identified. Therefore, we should have $\alpha\neq-\frac{1}{\theta}$ to ensure the non-identification of the conduct parameter.
A.5 Proof of Lemma 5
From the proof of Lemma 4, we know that $C_{i}$ and $D_{i}$ are independent of $\bm{X}^{d}$ under (13). Under the inverse demand function (89), we have
$\hat{C}_{i}(Q,\bm{X}^{d})=\frac{Q^{\alpha}r_{i}(\bm{X}^{d})+\theta^{*}\alpha Q^{\alpha}r_{i}(\bm{X}^{d})}{Q^{\alpha}r_{i}(\bm{X}^{d})+\theta\alpha Q^{\alpha}r_{i}(\bm{X}^{d})}=\frac{1+\theta^{*}\alpha}{1+\theta\alpha}.$ (110)
@@ A.5 Proof of Lemma 5
$\hat{D}_{i}(Q,\bm{X}^{d})$ $=\frac{\theta^{*}-\theta}{Q^{\alpha}+\theta\alpha Q^{\alpha}}\left[s^{\prime}(Q)Q^{\alpha}+Qs^{\prime\prime}(Q)Q^{\alpha}-s^{\prime}(Q)\alpha Q^{\alpha}\right]$ (111)
@@
The last equality holds because $Q^{e}$ satisfies the equilibrium condition under $(\theta,MC)$. Therefore, the two models are observationally equivalent, which implies that the non-identification holds.
@@ Appendix B Summary of Goldman and Uzawa (1964)
Then, Goldman and Uzawa (1964) specifies the functional form that a weak separable function should satisfy.
@@
where $\Phi(r^{1},\ldots,r^{K})$ is a function of $K$ variables and, for each $k$, $r^{k}(x^{k})$ is a function of subvector $x^{k}$ alone.